# Visual Sparse Steering: Improving Zero-shot Image Classification with Sparsity Guided Steering Vectors

### v1

@@ Results.
In Table 1, we present the results for VS2 ++ when external data is available under both RAG-enhanced settings where we have oracle and non-oracle cache data. When we have oracle positive and negative sets, we observe that VS2++ surpasses all other methods under almost every setting. Weighted RAG performs worse than the steering vector approaches (both CLIP and VS2++), suggesting that directionality is more informative than proximity. In the more realistic scenario where we do not have oracle positive and negative sets, we observe drops in accuracies. We hypothesize that this degradation is due to inaccuracies in pseudolabeling, resulting in the amplification of spurious directions in the SAE latent space. However, across the six configurations, VS2++ outperforms the CLIP Steering Vector in four cases and falls behind only in the ViT-B/32 runs for CIFAR-100 and Tiny-ImageNet (with CIFAR-100 essentially on par). These results suggest that there remains significant room for improvement in actively selecting which sparse features to steer, especially under noisy or weakly supervised retrieval scenarios. In Appendix I, we conduct ablation studies on the top-$N$ retrieved images that indicate that the utility of neighbor-based aggregation is dataset-dependent: general object recognition tasks benefit from larger $N$, while fine-grained classification may require more careful control of retrieval scope.
4.1.3 Fine-Grained Per-Class Accuracy Analysis
To understand where Visual Sparse Steering helps most, we compute per-class top-1 accuracy on CIFAR-100 with a ViT-B/32 backbone and report the ten largest improvements in Table 2. The results reveal a long-tailed pattern: VS2 lifts accuracy by up to 25 % in individual classes (e.g. lawn-mower $\rightarrow$ tractor, pine-tree $\rightarrow$ forest), while VS2++ reaches gains of 38 % when high-quality neighbours are available (e.g. bee $\rightarrow$ spider, lion $\rightarrow$ tiger, whale $\rightarrow$ flatfish). In all cases, the corrected predictions involve visually or taxonomically proximate categories, indicating that sparse steering primarily disambiguates fine-grained confusions rather than providing a uniform boost. This behaviour aligns with our qualitative findings in the Appendix A.2, where the top-$k$ SAE features capture subtle class-specific attributes (birds with gray upperpart, white underpart, white-colored throat and black eyes etc.) that are otherwise entangled in the original CLIP embedding. A more detailed list of per-class performance accuracies of VS2 and VS2++ is provided in Table 9 in the Appendix.
@@ 4.1.3 Fine-Grained Per-Class Accuracy Analysis
Table 2: Top-10 class gains on CIFAR-100. Numbers are top-1 accuracies; green = absolute gain over zero-shot (ZS).
Class ZS VS2 (+$\Delta$) Mislabel Why visually confusing tractor $55.0$ 80.0 $\uparrow$(+25.0) lawn_mower Both are wheeled agricultural machines; similar viewed side-on. forest $35.0$ 58.0 $\uparrow$(+23.0) pine_tree Dense tree canopies often dominated by tall conifers. man $53.0$ 74.0 $\uparrow$(+21.0) boy Male human silhouette; age difference subtle at CIFAR resolution. bus $51.0$ 68.0 $\uparrow$(+17.0) pickup_truck Rectangular vehicle profile with windows and wheels. snake $61.0$ 76.0 $\uparrow$(+15.0) worm Long, legless bodies with smooth textures. woman $57.0$ 71.0 $\uparrow$(+14.0) girl Female human figures; height/size cues are small. trout $18.0$ 32.0 $\uparrow$(+14.0) aquarium_fish Similar fish shape and reflective scales in water scenes. cockroach $49.0$ 62.0 $\uparrow$(+13.0) beetle Small, dark exoskeleton insects with segmented bodies. bridge $72.0$ 83.0 $\uparrow$(+11.0) sea Large spans photographed above wide blue water areas. rose $58.0$ 68.0 $\uparrow$(+10.0) poppy Red, multi-petal flowers with central dark core.
(a) VS2
Class ZS VS2++ (+$\Delta$) Mislabel Why visually confusing spider $53.0$ 91.0 $\uparrow$(+38.0) bee Small multi-legged insects; dark radial silhouettes. tiger $48.0$ 84.0 $\uparrow$(+36.0) lion Large orange/brown big cats; stripes vs. mane often unclear at low resolution. flatfish $49.0$ 85.0 $\uparrow$(+36.0) whale Flattened marine shapes against blue water backgrounds. possum $30.0$ 65.0 $\uparrow$(+35.0) hamster Small brown furry mammals with rounded bodies. lizard $39.0$ 74.0 $\uparrow$(+35.0) snake Slender reptiles; legs on lizard hard to spot in $32\times 32$ images. sweet_pepper $58.0$ 92.0 $\uparrow$(+34.0) orange Round orange-coloured produce with glossy skins. raccoon $49.0$ 83.0 $\uparrow$(+34.0) skunk Mid-sized nocturnal mammals with dark fur and contrasting markings. shrew $46.0$ 79.0 $\uparrow$(+33.0) skunk Small ground mammals; dark elongated silhouettes blur distinguishing cues. snail $53.0$ 86.0 $\uparrow$(+33.0) mushroom Brown dome (shell vs. cap) resting on the forest floor. beetle $49.0$ 82.0 $\uparrow$(+33.0) cockroach Dark, shiny exoskeletal insects with similar body shapes.
(b) VS2++
4.2 Prototype-Aware Sparse Steering Vectors
Our core methods, VS2 and its retrieval-augmented variant VS2++, enhance zero-shot classification by steering CLIP embeddings along directions identified by a top-$k$ Sparse Autoencoder (SAE). These directions correspond to latent concepts that, ideally, align with class-discriminative features. Steering in these directions upweights what the model has learned to be important during reconstruction. This raises a central hypothesis: the reconstruction task itself is to some extent sufficient to uncover features that are also relevant for classification. In other words, there is a meaningful overlap between features that are important for reconstructing the CLS token and those that are predictive for the downstream task. In this section, drawing inspiration from prototype theory rosch1973natural , we investigate whether incorporating prototype information during SAE training can better align the features important for reconstruction with those that are critical for downstream classification.
Oracle steering with known prototypes.
Following prototype theory, we collect for every class the ten images that ViT-B/32 CLIP classifies with the highest confidence; these serve as oracle prototypes. Given the true class label $y$, we build a prototype steering vector by averaging their latent sparse features. VS2 using oracle prototypes can lift CLIP to 97.5% on CIFAR-100, 91.04% on CUB-200, and 90.1% on Tiny-ImageNet. This confirms the existence of discriminative sparse directions. In Appendix E we further examine the discriminative ability (measured by orthogonality) of these steering vectors. Generally, these prototype vectors have a low cosine similarity, yet a non-negligible tail reveals strongly overlapping directions between visually or semantically close categories.
Prototype-aligned SAE (PASS).
@@ 5 Conclusion
We introduce Visual Sparse Steering VS2 and its retrieval-augmented variant VS2++, two test-time methods that steer vision foundation models using sparse features obtained from top-$k$ Sparse Autoencoders without requiring contrastive data. Specifically, VS2 consistently outperforms all baseline methods across all datasets and ViT backbones, surpassing zero-shot CLIP by 3.45% and 4.12% on CIFAR-100, 0.93% and 1.08% on CUB-200, and 1.50% and 1.84% on Tiny-ImageNet with ViT-B/32 and ViT-B/16 backbones, respectively. To control concept importance more precisely, we propose VS2++. With oracle positive/negative sets, VS2++ surpasses every baseline in nearly every configuration, achieving absolute top-1 gains over CLIP zero-shot of up to 21.44% on CIFAR-100, 7.08% on CUB-200, and 20.47% on Tiny-ImageNet. When those sets are noisy, its accuracy drops but still surpasses the standard Steering Vector in most cases, highlighting the need for better feature selection under weakly supervised retrieval. Interestingly, VS2 and VS2++ raise per-class accuracy by up to 25% and 38%, respectively, showing that sparse steering benefits specific classes by disambiguating visually or taxonomically proximate categories rather than providing a uniform boost. Finally, to further align the sparse features learned through the SAE reconstruction task with those important for downstream performance, we proposed Prototype-Aligned Sparse Steering (PASS). By adding a lightweight prototype-alignment loss during SAE training (labels used only in training, still test-time unsupervised) PASS consistently, though slightly, outperforms VS2, offering a marginal 6.12% gain over VS2 only on CIFAR-100 with ViT-B/32. Selecting appropriate prototypes or exploring multi-prototype guidance is an intriguing avenue for future work.
@@ Appendix A Decoding the Sparse Latent Space: Insights from SAEs
Our proposed methods, VS2 and VS2++, achieve significant classification performance gains, primarily due to the contribution of Sparse Autoencoders (SAEs). We hypothesize that SAEs identify meaningful sparse concepts, which in turn guide the steering mechanisms. To validate this assumption, we conduct both quantitative and qualitative evaluations to assess the significance of the learned concepts.
A.1 Quantitative Evaluation of Concept Significance.
To further evaluate the role of interpretable sparse features, we manipulate the top-$k$ most active latent dimensions extracted via Sparse Autoencoders. This experiment examines whether these features are critical for classification and how model predictions change under different modifications. We explore the following manipulation settings:
@@
Table 3 reports the zero-shot classification accuracy of $\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ using ViT-B/32 after applying these transformations to the Sparse Steering vector intervention. We observe that negating or zeroing the top-$k$ most important concepts consistently degrades performance across all datasets. This result confirms that the concepts learned by the Sparse Autoencoders are essential for classification.
@@
Modification CIFAR-100 CUB-200 Tiny-IN
$\text{CLIP}_{\text{ZS}}$ $61.07$ $38.68$ $56.64$
$\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ $62.69$ $37.30$ $39.49$
+ Zero-out ($\gamma=0$) $1.71$ $1.14$ $16.11$
+ Negate ($\gamma=-1$) $0.06$ $0.00$ $0.82$
A.2 Qualitative Evaluation of Concept Significance.
We qualitatively investigate the concepts learned in the sparse latent representations of the top-$k$ activations in the Sparse Autoencoder (SAE). Specifically, we assess the learned features by analyzing feature activations for each input and identifying the inputs that exhibit the highest activations for a given feature. Unlike mechanistic interpretability in the language domain, where an LLM can be used to assign semantic labels to a feature by summarizing its highly activated inputs, the vision domain lacks an equivalent automated labeling process.
To avoid reliance on human qualitative evaluation, we leverage annotated datasets where each image is associated with predefined attributes. For the qualitative evaluation of concept significance, we use the CUB dataset, which provides rich concept annotations for each image, enabling a structured assessment of the learned representations. Specifically, we investigate whether we can identify specific latent features with the highest concept coverage among their top-k most activated images. Concept coverage refers to how consistently a specific interpretable concept (e.g., an identifiable object category, attribute, or semantic idea) appears across a set of highly activated examples for a given SAE dimension. The intuition is that if a specific concept frequently emerges among the top-activating images for a particular dimension, that dimension is strongly associated with that concept.

### v2

@@ Results.
Table 1 reports VS2++ under oracle (upper bound) and non-oracle (CLIP pseudo-labeled) positive/negative sets. In the oracle setting, VS2++ achieves gains of up to 21.44% (CIFAR-100), 7.08% (CUB-200), and 20.47% (Tiny-ImageNet), demonstrating substantial headroom when feature selection is accurate. Weighted RAG is generally weaker than contrastive steering on CIFAR-100 and Tiny-ImageNet while remaining competitive on CUB-200, indicating that directional offsets are often more informative than proximity alone; VS2++ is competitive with CLIP-space steering but not uniformly best (e.g., Tiny-ImageNet with ViT-B/32). In the non-oracle setting, accuracy drops due to pseudo-label noise, yet VS2++ outperforms the CLIP steering vector in four of six configurations and does not exhibit negative transfer on CUB-200. Overall, these results motivate robust feature selection under weak supervision; Appendix I ablates the number of retrieved neighbors.
@@
Table 4: Top-5 class gains on CIFAR-100. Top-1 accuracy; green = gain over CLIP zero-shot.
Class ZS VS2 / VS2++ Mislabel Why visually confusing VS2 tractor $55.0$ 80.0 $\uparrow$(+25.0) lawn_mower wheeled farm machines; side view forest $35.0$ 58.0 $\uparrow$(+23.0) pine_tree dense conifer canopy man $53.0$ 74.0 $\uparrow$(+21.0) boy age cue weak at $32{\times}32$ bus $51.0$ 68.0 $\uparrow$(+17.0) pickup_truck boxy vehicle silhouette snake $61.0$ 76.0 $\uparrow$(+15.0) worm elongated legless body VS2++ spider $53.0$ 91.0 $\uparrow$(+38.0) bee small dark radial silhouette tiger $48.0$ 84.0 $\uparrow$(+36.0) lion stripes vs. mane unclear flatfish $49.0$ 85.0 $\uparrow$(+36.0) whale flat marine shape on blue possum $30.0$ 65.0 $\uparrow$(+35.0) hamster small brown furry mammal lizard $39.0$ 74.0 $\uparrow$(+35.0) snake legs often invisible
4.3 Fine-Grained Per-Class Accuracy Analysis
[…]
4.4 SAEs as a reliability diagnostic for safe visual sparse steering
Our main experiments assume access to a sparse autoencoder whose reconstructions of the vision encoder activations are sufficiently faithful on the target distribution. We approximate this by training top-$k$ SAEs on unlabeled in-domain training-split activations for each dataset, which allows us to isolate and investigate the effectiveness of sparsity-guided steering vectors in modulating the classification behavior of CLIP’s vision transformer.
Relaxation: a generalized SAE.
In practice, a single high-fidelity SAE that remains faithful across diverse domains is difficult to obtain. To approximate this setting, we train a generalized SAE on the union of the unlabeled training splits of CIFAR-100, CUB-200, and Tiny-ImageNet and then apply this same SAE across datasets. Table 5 reports top-1 accuracy together with Fraction of Variance Unexplained (FVU) for generalized steering. Larger FVU values indicate poorer reconstruction, and FVU >1>1 implies worse reconstruction than a mean predictor. We observe that generalized steering improves over the zero-shot baseline on CIFAR-100 and Tiny-ImageNet, where reconstruction error is low (FVU $\approx 0.21$ on CIFAR-100 and $\approx 0.44$ on Tiny-ImageNet), but degrades on CUB-200, where reconstruction error is substantially higher (FVU $\approx 1.93$). This pattern suggests that steering effectiveness is sensitive to SAE fidelity, motivating the use of reconstruction error as a reliability signal.
Table 5: Reconstruction and classification with a generalized SAE. We report top-1 accuracy (%) with corresponding reconstruction error FVU in parentheses.
CIFAR-100 CUB-200 Tiny-IN
Method B/32 B/16 B/32 B/16 B/32 B/16
Baseline 61.07 (–) 63.96 (–) 51.76 (–) 55.06 (–) 56.64 (–) 61.08 (–)
VS2 64.63 (0.216) 68.22 (0.208) 48.81 (1.931) 49.22 (1.930) 59.73 (0.448) 63.83 (0.437)
Table 6: FVU-calibrated fallback for VS2 on in-distribution and out-of-distribution (OOD) datasets using a generalized SAE. Each cell reports top-1 accuracy with coverage (fraction steered) in parentheses.
In Distribution Out of Distribution
Method CIFAR-100 CUB-200 Tiny-IN Aircraft Food101 Flower102 Caltech101 SUN EuroSAT
[…]
VS2 (In Domain SAE, ungated) 64.52 52.69 58.14 27.69 81.20 64.81 80.95 54.80 34.65
VS2 (Generalized SAE, ungated) 64.63 48.81 59.73 18.57 81.06 64.11 79.72 45.78 30.76
VS2 + FVU gate (q=0.90q{=}0.90) 64.27 (99.93%) 51.76 (0.0%) 59.29 (85.0%) 24.87 (0.00%) 80.72 (0.00%) 63.70 (0.11%) 78.44 (2.30%) 51.80 (0.11%) 30.39 (54.20%)
VS2 + FVU gate (q=0.95q{=}0.95) 64.62 (99.97%) 51.76 (0.04%) 58.44 (92.16%) 24.87 (0.00%) 80.71 (0.02%) 63.72 (1.32%) 78.52 (5.51%) 51.80 (0.49%) 31.39 (63.15%)
VS2 + FVU gate (q=0.99q{=}0.99) 64.63 (100%) 51.79 (0.62%) 58.65 (98.63%) 24.87 (0.09%) 80.70 (0.89%) 64.34 (22.72%) 78.76 (23.82%) 51.92 (6.97%) 31.57 (85.52%)
VS2 + FVU gate (q=0.995q{=}0.995) 64.63 (100%) 51.83 (1.54%) 58.64 (99.34%) 24.87 (0.21%) 80.69 (3.10%) 64.21 (40.62%) 78.78 (35.26%) 52.11 (15.28%) 31.65 (92.09%)
Out-of-distribution steering with FVU-gated fallback.
We next evaluate robustness under distribution shift by applying the generalized SAE to six unseen datasets: Aircraft (Maji et al., 2013), Food101 (Bossard et al., 2014), Flower102 (Nilsback and Zisserman, 2008), Caltech101 (Fei-Fei, 2004), SUN (Xiao et al., 2010), and EuroSAT (Helber et al., 2019). As a reference point, Table 6 shows that when an in-domain SAE is used, VS2 consistently improves over zero-shot CLIP across all six OOD datasets. However, ungated generalized steering can substantially degrade performance on some OOD datasets (e.g., Aircraft and SUN), consistent with unreliable sparse features when reconstruction quality is poor. We therefore use per-sample FVU as a reliability diagnostic: if a test sample’s FVU exceeds a threshold $\tau$, we skip steering and fall back to the baseline CLIP embedding; otherwise, we apply VS2.
To set $\tau$ without labels, we calibrate a single threshold on the unlabeled CIFAR-100+Tiny-ImageNet training union by taking the $q$-quantile of per-sample FVU scores, and then apply the same $\tau$ across all datasets. For FVU-gated rows in Table 6, we report top-1 accuracy together with coverage, the fraction of samples for which steering is applied. Overall, FVU gating acts as a conservative safety mechanism: on datasets where generalized steering is harmful, coverage drops and performance moves toward the baseline, while on in-distribution datasets with reliable reconstruction, coverage remains high and the gains of generalized steering are largely retained.
5 Discussion and Future Work
We introduced VS2, a lightweight and fully label-free adaptation method that steers a frozen vision encoder using steering vectors derived from sparse features learned by top-$k$ Sparse Autoencoders (SAEs). Unlike conventional steering vectors that rely on curated contrastive anchors, VS2 constructs a per-instance direction directly from the SAE’s sparse support: it amplifies the active latents, projects the induced change back to embedding space, and applies the resulting offset at inference time. Across CIFAR-100, CUB-200, and Tiny-ImageNet and two CLIP backbones, VS2 consistently improves zero-shot accuracy while remaining forward-only and adding negligible compute overhead. More broadly, our results support the view that SAE latents can serve as an actionable intervention space for improving downstream behavior, beyond post-hoc interpretability.
A key practical constraint is SAE fidelity. When reconstruction is poor, the resulting sparse features can be unreliable and steering may become harmful. A central advantage of VS2 is that reconstruction error provides a model-internal reliability signal that is typically absent from standard steering-vector pipelines and many test-time adaptation methods. This enables a conservative, label-free fallback: when reconstruction error is high, we revert to the zero-shot prediction, preventing negative steering under distribution shift. An important future direction is to adapt the SAE per instance at test time via lightweight self-reconstruction to reduce error, improve the salient features, and therefore steering effectiveness.
Finally, our upper-bound analyses indicate substantial headroom when task-relevant sparse features can be selected reliably. This highlights a central gap: the SAE objective surfaces features that are salient for reconstruction, but the features most useful for downstream classification can differ and are task-dependent. In our current setup, the same SAE features are used independent of the task, and we uniformly amplify the active features. Bridging reconstruction saliency and task saliency motivates future work on task-aligned sparse steering, i.e., selecting and amplifying the subset of SAE features that most directly support the target decision. Appendix K takes an initial step in this direction with Prototype-Alignment Sparse Steering, inspired by prototype theory (Rosch, 1973), suggesting that incorporating weak or indirect task structure during representation learning can better align sparse features with downstream needs.
@@ 6 Conclusion
We introduced Visual Sparse Steering (VS2), a lightweight, label-free domain adaptation method that steers vision foundation models using sparse features learned by top-$k$ Sparse Autoencoders, without updating model weights or requiring contrastive anchor examples. Across CIFAR-100, CUB-200, and Tiny-ImageNet and two CLIP backbones (ViT-B/32, ViT-B/16), VS2 improves over zero-shot CLIP by 3.45%–4.12%, 0.93%–1.08%, and 1.50%–1.84%, respectively, while remaining forward-only at inference. To study the potential benefits of selective feature amplification, we presented VS2++ as a retrieval-based upper-bound analysis that uses retrieved neighbors to form positive/negative groups and construct a contrastive steering direction in SAE feature space. With oracle positives/negatives, VS2++ yields absolute top-1 gains over CLIP zero-shot of up to 21.44% on CIFAR-100, 7.08% on CUB-200, and 20.47% on Tiny-ImageNet, illustrating substantial headroom when feature selection is accurate. When the retrieved sets are noisy, performance can degrade, highlighting the need for more robust feature selection under weak supervision. We additionally observe that VS2 and VS2++ can disproportionately benefit visually or taxonomically proximate classes, improving per-class accuracy by up to 25% and 38%, respectively. Finally, VS2 provides a reconstruction-error-based reliability diagnostic that can disable steering and fall back to the baseline under a strict threshold, reducing the risk of harmful interventions at the cost of occasionally foregoing gains.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Decoding the Sparse Latent Space: Insights from SAEs
Our proposed methods achieve significant classification performance gains, primarily due to the contribution of Sparse Autoencoders (SAEs). We hypothesize that SAEs identify meaningful sparse features, which in turn guide the steering mechanisms. To validate this assumption, we conduct both quantitative and qualitative evaluations to assess the significance of the learned features.
A.1 Quantitative Evaluation of Feature Significance
To evaluate the role of sparse features, we manipulate the top-$k$ most active latent features extracted via Sparse Autoencoders. This experiment examines whether these features are critical for classification and how model predictions change under different modifications. We explore the following manipulation settings:
@@ A.1 Quantitative Evaluation of Feature Significance
Table 7 reports the zero-shot classification accuracy of $\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ using ViT-B/32 after applying these transformations to the Sparse Steering vector intervention. We observe that negating or zeroing the top-$k$ most important features consistently degrades performance across all datasets. This result confirms that the features learned by the Sparse Autoencoders are essential for classification.
@@
Modification CIFAR-100 Tiny-IN
$\text{CLIP}_{\text{ZS}}$ $61.07$ $56.64$
$\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ $62.69$ $39.49$
+ Zero-out ($\gamma=0$) $1.71$ $16.11$
+ Negate ($\gamma=-1$) $0.06$ $0.82$
A.2 Qualitative Evaluation of Features Significance
We qualitatively investigate the features learned in the sparse latent representations of the top-$k$ activations in the Sparse Autoencoder (SAE). Specifically, we assess the learned features by analyzing feature activations for each input and identifying the inputs that exhibit the highest activations for a given feature. Unlike mechanistic interpretability in the language domain, where an LLM can be used to assign semantic labels to a feature by summarizing its highly activated inputs, the vision domain lacks an equivalent automated labeling process.
To avoid reliance on human qualitative evaluation, we leverage annotated datasets where each image is associated with predefined attributes. For the qualitative evaluation of feature significance, we use the CUB dataset, which provides rich concept annotations for each image, enabling a structured assessment of the learned representations. Specifically, we investigate whether we can identify specific latent features with the highest concept coverage among their top-k most activated images. Concept coverage refers to how consistently a specific interpretable concept (e.g., an identifiable object category, attribute, or semantic idea) appears across a set of highly activated examples for a given SAE dimension. The intuition is that if a specific concept frequently emerges among the top-activating images for a particular feature, that feature is strongly associated with that concept.
@@ A.2 Qualitative Evaluation of Features Significance
Table 8: Top-10 most activated classes for SAE Feature 511 (left) and Feature 3067 (right). Feature 511 strongly aligns with gull-like classes, while Feature 3067 captures a cross-class head/throat attribute.
Class (511) Activation Class (3067) Activation 066.Western_Gull 96.7% (29/30) 082.Ringed_Kingfisher 60.0% (18/30) 060.Glaucous_winged_Gull 96.6% (28/29) 083.White_breasted_Kingfisher 40.0% (12/30) 061.Heermann_Gull 93.3% (28/30) 079.Belted_Kingfisher 30.0% (9/30) 059.California_Gull 90.0% (27/30) 041.Scissor_tailed_Flycatcher 23.3% (7/30) 062.Herring_Gull 83.3% (25/30) 080.Green_Kingfisher 16.7% (5/30) 063.Ivory_Gull 76.7% (23/30) 002.Laysan_Albatross 13.3% (4/30) 087.Mallard 76.7% (23/30) 025.Pelagic_Cormorant 13.3% (4/30) 147.Least_Tern 76.7% (23/30) 025.Pelagic_Cormorant 13.3% (4/30) 064.Ring_billed_Gull 70.0% (21/30) 044.Frigatebird 10.0% (3/30) 084.Red_legged_Kittiwake 73.9% (17/23) 049.Boat_tailed_Grackle 10.0% (3/30)
Class-Conditional Feature Activation.
To provide additional evidence that SAE dimensions capture discriminative visual concepts, we report the class-conditional activation frequencies of two example features: 511 and 3067. Specifically, Table 8 shows the percentage of samples in the top-10 activating classes where each feature is present.
What we observe is that feature 511 fires very often in gull-like classes. It also fires Larus-type seabirds, i.e., terns, kittiwakes. We observe that it also fires moderately on similar seabirds (Horned Puffin: 53.3%, White Pelican: 40%, Brown Pelican: 16.7%, Red-breasted Merganser: 10%). Additionally, we observe that Feature 511 activates near 0% on unrelated classes such as buntings, warblers, and hummingbirds. In contrast, Feature 3067 captures a cross-category visual attribute rather than a taxonomic grouping. Specifically, it consistently activates on species that share a distinctive head-and-throat pattern (e.g., white throat with darker eye region), even when those species are taxonomically unrelated. This indicates that the SAE possibly learns attribute-level discriminative concepts, complementing the class-specific features exemplified by Feature 511.
Discussion.
We note that our goal is not to claim human interpretability of SAE features. As shown in recent mechanistic interpretability work (Chanin et al., ), human-aligned features are often split across multiple SAE components (feature absorption), a known behavior in sparse dictionary learning. Nonetheless, our contribution lies in demonstrating that SAEs can go beyond interpretability, serving as effective mechanisms for performance improvement via sparse steering. While some features are interpretable, our main result is that sparse features enable controllable improvements in downstream classification tasks.